OpenStudy (anonymous):

Use the slope and y-intercept to graph each equation. -2(3x-4)+y=0 What is the slope and y-intercept and how do i find it?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

um first distribute?

OpenStudy (skullpatrol):

@Lizzychey The first thing you need to know is what form do you want to put the given equation into?

OpenStudy (anonymous):

y=mx+b

OpenStudy (skullpatrol):

Very good. :) This is called the slope intercept form where m is the slope and b is the y-intercept.

OpenStudy (anonymous):

right but how am i suppose to find them when its -2(3x-4)+y=0 its nowhere near y=mx+b

OpenStudy (jim766):

distribute the -2 (3x -4) + y = 0

OpenStudy (anonymous):

-6x+8+y=0

OpenStudy (skullpatrol):

subtract 8 from both sides

OpenStudy (anonymous):

why not -6x?

OpenStudy (skullpatrol):

not yet

OpenStudy (anonymous):

ok so -6x+y=-8

OpenStudy (skullpatrol):

Good, now add 6x to both sides.

OpenStudy (anonymous):

then add -6x to both sides which gives me y=6x-8

OpenStudy (skullpatrol):

6x +(-6x) = 0

OpenStudy (anonymous):

ok so now its in y=mx+b form so the slope is 6 and the y-intercept is -8?

OpenStudy (anonymous):

so the 2 points are (0,-8), (1,-2) ?

OpenStudy (skullpatrol):

Correct :)

OpenStudy (anonymous):

Yay!! Thank you! :)

OpenStudy (skullpatrol):

np :)